# Semiregular abstract polyhedra with trivial facet stabilizer

### v1

@@ 4 Voltage graphs
To define voltage graphs first we need to think of graphs in a way that is not completely standard. For this paper, a graph $\mathcal{X}$ consists of a set of vertices $V(\mathcal{X})$, a set of darts $D(\mathcal{X})$, a function $I:D(\mathcal{X})\to V(\mathcal{X})$ and an involution $(\cdot)^{-1}:D(\mathcal{X})\to D(\mathcal{X})$. We call $I(d)$ the initial vertex of $d$ and $d^{-1}$ the inverse of $d$. An edge of $\mathcal{X}$ is just a set of the form $\{d,d^{-1}\}$. A semi-edge is a dart $d$ satisfying $d=d^{-1}$. If $d$ is a semi-edge, we abuse language and call the edge $\{d\}$ a semi-edge as well. We think of the edge $e=\{d,d^{-1}\}$ as a line joining the vertices $I(d)$ and $T(d):=I(d^{-1})$ (called the final vertex of $d$), and we think of its darts as the two (or one) possible orientations. A path $W$ on $\mathcal{X}$ is a finite sequence $d_{1}d_{2}\cdots d_{\ell}$ of darts in $\mathcal{X}$ such that $T(d_{i})=I(d_{i+1})$ for all $i\in\{0,1,\ldots,\ell-1\}$. The vertex $I(d_{1})$ is the initial vertex of $W$ while $T(d_{\ell})$ is its final vertex. The number $\ell$ is called the length of $W$. We refer the reader to [9] for details.
@@
Given a voltage graph $(\mathcal{X},\Gamma,\xi)$ we can construct a graph $\mathcal{X}^{\xi}$ where $\Gamma$ acts by automorphisms. This action is regular on both the vertices and darts of $\mathcal{X}$ and the quotient $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$. To define $\mathcal{X}^{\xi}$ we make $V(\mathcal{X}^{\xi})=V(\mathcal{X})\times\Gamma$ and $D(\mathcal{X}^{\xi})=D(\mathcal{X})\times\Gamma$. Then we define the initial vertex of the dart $(d,\gamma)$ to be $I(d,\gamma)=(I(d),\gamma)$; and the inverse of $(d,\gamma)$ to be $(d,\gamma)^{-1}=(d^{-1},\xi(d)\gamma)$. In other words, if there is a dart $d$ from $x$ to $y$ in $\mathcal{X}$, then there is a dart $(d,\gamma)$ from $(x,\gamma)$ to $(y,\xi(d)\gamma)$ in $\mathcal{X}^{\xi}$. If $\mathcal{X}$ has a coloring (of its vertices, darts and/or edges) we color $(x,\gamma)$ with the same color as $x$.
One can verify that the $\Gamma$-action $(x,\gamma)\tau=(x,\gamma\tau)$ (where $x$ might be either a vertex or a dart) on $\mathcal{X}^{\xi}$ is by (color-preserving) automorphisms, regular and that two elements (darts or vertices) are in the same $\Gamma$-orbit if and only if they have the same first coordinate. Therefore, $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$. This gives a natural embedding of $\Gamma$ into the automorphism group of $\mathcal{X}^{\xi}$.
It is convinient to work with voltage graphs where $\mathcal{X}$ has a spanning tree $T$ satisfying that all its darts have trivial voltage. In this case, each connected component of $\mathcal{X}^{\xi}$ has $k|V(\mathcal{X})|$ vertices, where $k=|\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle|$. To be more precise, the connected component of the vertex $(x,\gamma)$ has vertex set $V(\mathcal{X})\times\gamma\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle$.
@@
  • Whenever $|i-j|>1$, the voltage of any alternating path of length 4 with colors $i$ and $j$ is trivial.
In [7, Lemma 3.1] it is proven that if $(\mathcal{X},\Gamma,\xi)$ is a voltage premaniplex, then $\mathcal{X}^{\xi}$ is a maniplex. As mentioned before, $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$, so if one can prove that every automorphism of $\mathcal{X}^{\xi}$ can be interpreted as an element of $\Gamma$, then $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
@@
[9, Theorem 4.1] Let $(X,\Gamma,\xi)$ be a voltage graph such that $\mathcal{X}^{\xi}$ is connected. Let $x$ be a vertex of $\mathcal{X}$ and $\tau$ an automorphism of $X$. Then $\tau$ lifts if and only if the function that maps $\xi(W)$ to $\xi(W\tau)$ for every closed path $W$ based at $x$ is well defined. In this case, this function is an automorphism of $\Gamma$.
@@
[1, Corollary 5.2] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage premaniplex. If $\mathcal{X}$ is regular, then every automorphism of $\mathcal{X}^{\xi}$ projects.
@@ 5 Regular symmetry type graphs of semiregular facet-transitive polyhedra with trivial facet stabilizer
In this section we will characterize those regular premaniplexes that are symmetry type graps of semiregular, facet-trasnitive polyhedra with trivial facet stabilizer. Recall that, since all abstract polygons are regular, an abstract polyhedron is semiregular if and only if it is vertex-transitive, so we are looking for polyhedra that are both vertex- and facet-transitive.
Let us ask: What could be the symmetry type graph $\mathcal{X}$ of a facet-transitive semiregular ($n+1)$-polytope $\mathcal{P}$ with trivial facet stabilizer? As established in Section 3, $\mathcal{X}_{\overline{n}}$ must be connected for $\mathcal{P}$ to be facet-transitive, and it must be isomorphic to a facet of $\mathcal{P}$ to ensure that facet-stabilizers are trivial in $\mathcal{P}$. Therefore, to construct $\mathcal{X}$ we may start with an $n$-polytope $\mathcal{K}$ (which will be isomorphic to a facet of $\mathcal{P}$), and then add edges of color $n$ in such a way that:
@@
So, for $n=3$ we are looking for a 3-premaniplex $\mathcal{X}$ satisfying that $\mathcal{X}_{\overline{2}}$ is a polygon, and that $\mathcal{X}_{\overline{0}}$ is connected. In order to be able to use 4.2, we will look for examples where $\mathcal{X}$ is also regular.
@@
Start with an $n$-gon, that is, a 2-maniplex with $2n$ flags forming an alternating cycle with colors 0 and 1. We want to add edges of color 2 that ‘‘commute’’ with the edges of color 0 in such a way that our conditions are satisfied. First notice that if there is a semi-edge of color 2, because of regularity, all edges of color 2 would be semi-edges. In this case $\mathcal{X}_{\overline{0}}$ would have $n$ connected components (each consisting of an edge of color 1 and two semi-edges of color 2), and this contradicts our hypothesis that $\mathcal{X}_{\overline{0}}$ is connected. Therefore, there cannot be any semi-edge in $\mathcal{X}$. Since the alternating paths of length 4 with colors 0 and 2 are closed, using regularity we get that either they are all 4-cycles or they are all 2-cycles that are travelled twice. If these paths are 2-cycles, it means that every edge of color 2 is parallel to an edge of color 0. If $\mathcal{X}$ is an $n$-gon with aditional edges of color 2 parallel to its edges of color 0, then $\mathcal{X}_{\overline{0}}$ is an alternating cycle of length $2n$ with colors 1 and 2, which is connected. Therefore, we have found our first family of premaniplexes with the desired properties: $n$-gons with edges of color 2 connecting pairs of 0-adjacent flags (see Figure 3). We will call this family 1.
@@
Now suppose that $\mathcal{X}$ satisfies our conditions and it has a 4-cycle with flags $\Phi$, $\Phi^{0}$, $\Phi^{02}$ and $\Phi^{2}$. There is an automorphism $\tau$ of $\mathcal{X}$ that maps $\Phi$ to $\Phi^{0}$ and $\Phi^{2}$ to $\Phi^{02}$. This means that $\tau$ fixes both the 1-face $e$ that connects $\Phi$ to $\Phi^{0}$ and $e^{\prime}$ that connects $\Phi^{2}$ to $\Phi^{02}$. The only way two different edges of a polygon are fixed by the same automorphism is if they are opposites in a polygon with an even number of sides. So we know that $n$ must be even and that 2-adjacent flags lie on opposite edges. The edge opposite to $e$ has the flags $(r_{1}r_{0})^{n/2}\Phi$ and $r_{0}(r_{1}r_{0})^{n/2}\Phi$, and this edge must be $e^{\prime}$. So, $\Phi^{2}$ must be one of these two flags. If $\Phi^{2}=(r_{1}r_{0})^{n/2}\Phi$ we get an alternating 4-cycle of colors 1 and 2 with flags $\Phi$, $\Phi^{2}$, $\Phi^{21}$ and $\Phi^{1}$. This means that for $n\geq 3$, $\mathcal{X}_{\overline{0}}$ cannot be connected. For $n=2$, however, we have found one sporadic example $\mathcal{X}_{s}$: a digon with edges of color 2 between opposite flags (see Figure 4).
@@
Let the edges (1-faces) of $\mathcal{X}$ be called $e_{0},e_{1},\ldots,e_{n-1}$ where $e_{0}$ is the edge of the flag $\Phi$ and the other edges are named cyclically in clockwise order. We will call the flags of the form $(r_{1}r_{0})^{k}\Phi$ left flags, and the flags of the form $r_{0}(r_{1}r_{0})^{k}\Phi$ right flags. Notice that the alternating path of length 4 with flags $\Phi$, $\Phi^{2}$, $\Phi^{21}$ and $\Phi^{212}$ has:
@@
We want this path to go through every flag before going back to $\Phi$. The only way to make sure that all these flags are different is if $n/2$ is even, so that each 4 steps have a left flag and a right flag of an even edge and of an odd edge. Therefore, our second family of premaniplexes with the desired conditions are $n$-gons where $n$ is divisible by 4, and the edges of color 2 connect each flag to the 0-adjacent flag of its opposite flag (see Figure 5). We will call this family 2.
@@
As discussed before, a semiregular facet-transitive 3-maniplex with trivial facet stabilizer and regular symmetry type graph must have a premaniplex $\mathcal{X}$ satisfying the conditions of 5.1 as its symmetry type graph. However, it is impossible to have a polyhedron $\mathcal{P}$ with trivial facet stabilizer and with $\mathcal{X}_{s}$ as its symmetry type graph: this is because the facets of $\mathcal{P}$ would have to be digons and all the polyhedra with only digonal faces are regular. So we get the following ressult:
@@
Let $\mathcal{M}$ be a semiregular facet-transitive 3 maniplex with trivial facet stabilizer. Let $\mathcal{X}$ be the symmetry type graph of $\mathcal{M}$. Then $\mathcal{X}$ is isomorphic to a premaniplex in either family 1 or family 2.
@@ 6 Constructing polyhedra with symmetry type graph in family 1
Now we want to assign voltages to the premaniplexes of the previous section to get examples of maniplexes and polytopes with these symmetry types. First, let $\mathcal{X}$ be the premaniplex of family 1 with $n$ sides. We want to find a voltage assignment $\xi$ on $\mathcal{X}$ such that $\mathcal{X}^{\xi}$ has symmetry type $\mathcal{X}$ and $n$-gonal faces. To ensure that $\mathcal{X}^{\xi}$ has $n$-gonal faces it is enough to assign trivial voltage to all the edges of color different than 2. Now let $\alpha_{0},\alpha_{1},\ldots,\alpha_{n-1}$ be the voltages of the edges of color 2 in cyclical order. To make sure that $\mathcal{X}^{\xi}$ is a maniplex we only need to make sure that $\alpha_{i}$ has order exactly 2 for every $i$. And finally, to make sure that $\mathcal{X}^{\xi}$ has symmetry type $\mathcal{X}$, we only need to make sure that no automorphism of $\mathcal{X}$ induces an automorphism of our voltage group $\Gamma=\langle\alpha_{0},\ldots,\alpha_{n-1}\rangle$. One way to do this is using the automorphism group of a regular $n$-polytope $\mathcal{P}$ as $\Gamma$ and making $\alpha_{i}=\rho_{i}$, where $\rho_{i}$ is the distinguished generator of $\Gamma(\mathcal{P})$ that maps a base flag $\Phi$ of $\mathcal{P}$ to its $i$-adjacent flag $\Phi^{i}$. If $\mathcal{P}$ is non-degenerate, then the only possible non-trivial permutation of distinguished generators that can induce an automorphism of $\Gamma(\mathcal{P})$ is the one mapping $\rho_{i}$ to $\rho_{n-1-i}$ for every $i$. This permutation corresponds to some reflection of $\mathcal{X}$ and it induces an automorphism of the group if and only if $\mathcal{P}$ is self-dual. So, if $\mathcal{P}$ is not self-dual and non-degenerate, $\mathcal{X}^{\xi}$ is a maniplex with symmetry type $\mathcal{X}$. This voltage assignment has been ilustrated in Figure 6, where the label of an edge of the polygon corresponds to the voltage of the darts of color 2 that start on that edge.
@@
[[1, Proposition 3.14]] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $(r+1)$-premaniplex. Suppose $\mathcal{X}_{\overline{r}}$ is connected, $\xi(d)=1$ for every dart $d$ of color smaller than $r$ and that $\mathcal{X}^{\xi}$ is a maniplex. Then $\mathcal{X}^{\xi}$ is a polytope if and only if $\mathcal{X}_{\overline{r}}$ is a polytope, and whenever there is a path $W$ with voltage 1 and colors in $[m,r]$, there is also a path $W^{\prime}$ with the same endpoints as $W$ that uses only colors in $[m,r-1]$
In our case $r=2$, so the only interesting case is when $m=1$. In other words, we want to make sure that if two flags are connected by a path with colors 1 and 2 and voltage 1, then they are either the same flag or 1-adjacent flags.
@@
This condition can be interpreted in terms of the Petrie polygon of $\mathcal{P}$. Typically, a Petrie polygon of a polytope is defined as a cyclical sequence of edges (1-faces) such that whenever $12\leq i\leq n-$ 1, any $i$ consecutive edges share an $i$-face but any $i+1$ consecutive edges do not. According to this definition some degenerate polytopes and some non-polytopal maniplexes would not have Petrie polygons, so the following alternative is prefered: We will define a Petrie polygon as a flag orbit under the action of the monodromy $\omega_{\pi}:=r_{n-1}r_{n-2}\cdots r_{0}$.The cardinallity $\ell$ of a Petrie polygon $P$ is called its length. If we look at the sequence of edges $\{(\omega_{\pi}^{i}\Phi)_{1}\}_{i=0}^{\ell}$ on a non-degenerate polytope, we would see that it satisfies the usual definition of Petrie polygon. Actually, in [18] the sequence $\{(\omega_{\pi}^{i}\Phi)_{1}\}_{i=0}^{\ell}$ is called a Petrie scheme and its sequence of edges is called a Petrie polygon, but in this article we do not make that distinction. Notice that the edge $(\omega_{\pi}^{i}\Phi)_{1}$ is incident to the vertices $(\omega_{\pi}^{i}\Phi)_{0}$ and $(\omega_{\pi}^{i+1}\Phi)_{0}$. In this definition we make sure that every flag is in exactly one Petrie polygon. If any two different flags in $P$ have different vertices (0-faces), we will say that $P$ is simple.
The following lemma will be quite useful:

### v2

@@ 4 Voltage graphs
Given a voltage graph $(\mathcal{X},\Gamma,\xi)$ we can construct a graph $\mathcal{X}^{\xi}$ where $\Gamma$ acts by automorphisms. This action is semiregular on both the vertices and darts of $\mathcal{X}^{\xi}$ and the quotient $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$. To define $\mathcal{X}^{\xi}$ we make $V(\mathcal{X}^{\xi})=V(\mathcal{X})\times\Gamma$ and $D(\mathcal{X}^{\xi})=D(\mathcal{X})\times\Gamma$. Then we define the initial vertex of the dart $(d,\gamma)$ to be $I(d,\gamma)=(I(d),\gamma)$; and the inverse of $(d,\gamma)$ to be $(d,\gamma)^{-1}=(d^{-1},\xi(d)\gamma)$. In other words, if there is a dart $d$ from $x$ to $y$ in $\mathcal{X}$, then for each $\gamma\in\Gamma$ there is a dart $(d,\gamma)$ from $(x,\gamma)$ to $(y,\xi(d)\gamma)$ in $\mathcal{X}^{\xi}$. If $\mathcal{X}$ has a coloring (of its vertices, darts and/or edges) we color $(x,\gamma)$ with the same color as $x$.
One can verify that the $\Gamma$-action $(x,\gamma)\tau=(x,\gamma\tau)$ (where $x$ might be either a vertex or a dart) on $\mathcal{X}^{\xi}$ is by (color-preserving) automorphisms, semiregular and that two elements (darts or vertices) are in the same $\Gamma$-orbit if and only if they have the same first coordinate. Therefore, $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$. This gives a natural embedding of $\Gamma$ into the automorphism group of $\mathcal{X}^{\xi}$.
It is convenient to work with voltage graphs where $\mathcal{X}$ has a spanning tree $T$ satisfying that all its darts have trivial voltage. In this case, each connected component of $\mathcal{X}^{\xi}$ has $k|V(\mathcal{X})|$ vertices, where $k=|\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle|$. To be more precise, the connected component of the vertex $(x,\gamma)$ has vertex set $V(\mathcal{X})\times\gamma\langle\xi(d):d\in D(\mathcal{X})\setminus D(T)\rangle$.
@@
  • Whenever |ij|>1|i-j|>1, the voltage of any alternating path of length 4 with colors $i$ and $j$ is trivial.
In [10, Lemma 3.1] it is proven that if $(\mathcal{X},\Gamma,\xi)$ is a voltage premaniplex, then $\mathcal{X}^{\xi}$ is a maniplex. As mentioned before, $\mathcal{X}^{\xi}/\Gamma$ is isomorphic to $\mathcal{X}$, so if one can prove that every automorphism of $\mathcal{X}^{\xi}$ can be interpreted as an element of $\Gamma$, then $\mathcal{X}$ is the symmetry type graph of $\mathcal{X}^{\xi}$.
@@
[12, Theorem 4.1] Let $(X,\Gamma,\xi)$ be a voltage graph such that $\mathcal{X}^{\xi}$ is connected. Let $x$ be a vertex of $\mathcal{X}$ and $\tau$ an automorphism of $\mathcal{X}$. Then $\tau$ lifts if and only if the function that maps $\xi(W)$ to $\xi(W\tau)$ for every closed path $W$ based at $x$ is well defined. In this case, this function is an automorphism of $\Gamma$.
@@
[3, Corollary 5.2] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage premaniplex. If $\mathcal{X}$ is regular, then every automorphism of $\mathcal{X}^{\xi}$ projects.
@@ 5 Regular symmetry type graphs of semiregular facet-transitive polyhedra with trivial facet stabilizer
In this section we will characterize those regular premaniplexes that are symmetry type graphs of semiregular, facet-transitive polyhedra with trivial facet stabilizer. Recall that, since all abstract polygons are regular, an abstract polyhedron is semiregular if and only if it is vertex-transitive, so we are looking for polyhedra that are both vertex- and facet-transitive.
Let us ask: What could be the symmetry type graph $\mathcal{X}$ of a facet-transitive semiregular ($n+1)$-polytope $\mathcal{P}$ with trivial facet stabilizer? As established in Section˜3, $\mathcal{X}_{\overline{n}}$ must be connected for $\mathcal{P}$ to be facet-transitive, and it must be isomorphic to a facet of $\mathcal{P}$ to ensure that facet-stabilizers are trivial in $\mathcal{P}$. Therefore, to construct $\mathcal{X}$ we may start with a regular $n$-polytope $\mathcal{K}$ (which will be isomorphic to a facet of $\mathcal{P}$), and then add edges of color $n$ in such a way that:
@@
So, for $n=3$ we are looking for a 3-premaniplex $\mathcal{X}$ satisfying that $\mathcal{X}_{\overline{2}}$ is a polygon, and that $\mathcal{X}_{\overline{0}}$ is connected. In order to be able to use ˜4.2, we will look for examples where $\mathcal{X}$ is also regular.
@@
Start with an $n$-gon, that is, a 2-maniplex with $2n$ flags forming an alternating cycle with colors 0 and 1. We want to add edges of color 2 that ‘‘commute’’ with the edges of color 0 in such a way that our conditions are satisfied. First notice that if there is a semi-edge of color 2, because of regularity, all edges of color 2 would be semi-edges. In this case $\mathcal{X}_{\overline{0}}$ would have $n$ connected components (each consisting of an edge of color 1 and two semi-edges of color 2), and this contradicts our hypothesis that $\mathcal{X}_{\overline{0}}$ is connected. Therefore, there cannot be any semi-edge in $\mathcal{X}$. Since the alternating paths of length 4 with colors 0 and 2 are closed, using regularity we get that either they are all 4-cycles or they are all 2-cycles that are traveled twice. If these paths are 2-cycles, it means that every edge of color 2 is parallel to an edge of color 0. If $\mathcal{X}$ is an $n$-gon with additional edges of color 2 parallel to its edges of color 0, then $\mathcal{X}_{\overline{0}}$ is an alternating cycle of length $2n$ with colors 1 and 2, which is connected. Therefore, we have found our first family of premaniplexes with the desired properties: $n$-gons with edges of color 2 connecting pairs of 0-adjacent flags (see Figure˜3). We will call this family 1.
@@
Now suppose that $\mathcal{X}$ satisfies our conditions and it has a 4-cycle with flags $\Phi$, $\Phi^{0}$, $\Phi^{02}$ and $\Phi^{2}$. There is an automorphism $\tau$ of $\mathcal{X}$ that maps $\Phi$ to $\Phi^{0}$ and $\Phi^{2}$ to $\Phi^{02}$. This means that $\tau$ fixes both the 1-face $e$ that consists of $\Phi$ and $\Phi^{0}$ and $e^{\prime}$ that consists of $\Phi^{2}$ and $\Phi^{02}$. The only way two different edges of a polygon are fixed by the same non-trivial automorphism is if they are opposites in a polygon with an even number of sides. So we know that $n$ must be even and that 2-adjacent flags lie on opposite edges. The edge opposite to $e$ has the flags $(r_{1}r_{0})^{n/2}\Phi$ and $r_{0}(r_{1}r_{0})^{n/2}\Phi$, and this edge must be $e^{\prime}$. So, $\Phi^{2}$ must be one of these two flags. If $\Phi^{2}=(r_{1}r_{0})^{n/2}\Phi$ we get an alternating 4-cycle of colors 1 and 2 with flags $\Phi$, $\Phi^{2}$, $\Phi^{21}$ and $\Phi^{1}$. This means that for $n\geq 3$, $\mathcal{X}_{\overline{0}}$ cannot be connected. For $n=2$, however, we have found one sporadic example $\mathcal{X}_{s}$: a digon with edges of color 2 between opposite flags (see Figure˜4).
@@
Let the edges (1-faces) of $\mathcal{X}$ be called $e_{0},e_{1},\ldots,e_{n-1}$ where $e_{k}$ is the edge that contains the flag $(r_{1}r_{0})^{k}\Phi$. We will call the flags of the form $(r_{1}r_{0})^{k}\Phi$ left flags, and the flags of the form $r_{0}(r_{1}r_{0})^{k}\Phi$ right flags. Notice that the alternating path of length 4 with flags $\Phi$, $\Phi^{2}$, $\Phi^{21}$ and $\Phi^{212}$ has:
@@
We want this path to go through every flag before going back to $\Phi$. The only way to make sure that all these flags are different is if $n/2$ is even, so that every 4 steps have a left flag and a right flag of an even edge and of an odd edge. Therefore, our second family of premaniplexes with the desired conditions are $n$-gons where $n$ is divisible by 4, and the edges of color 2 connect each flag to the 0-adjacent flag of its opposite flag (see Figure˜5). We will call this family 2.
@@
As discussed before, a semiregular facet-transitive 3-maniplex with trivial facet stabilizer and regular symmetry type graph must have a premaniplex $\mathcal{X}$ satisfying the conditions of ˜5.1 as its symmetry type graph. However, we will see that it is impossible for a 3-maniplex $\mathcal{M}$ with trivial facet stabilizer to have $\mathcal{X}_{s}$ as its symmetry type graph.
The $k$-gonal hosohedron ($k\in\mathbb{N}\cup\{\infty\}$, $k\geq 2$), denoted by $\{2,k\}$ is the regular polyhedron consisting of two vertices joined by $k$ edges forming $k$ digons. The $2k$-gonal hemi-hosohedron ($k\geq 1$ finite), denoted by $\{2,2k\}/2$, is the quotient of the $2k$-gonal hosohedron by identifying opposite flags. One might think of $\{2,2k\}/2$ as a map on the projective plane with a single vertex connected to itself by $k$ edges that touch it on opposite sides. In particular $\mathcal{X}_{s}=\{2,2\}/2$.
Lemma 5.2.
Let $\mathcal{M}$ be a 3-maniplex with only digonal faces. Then $\mathcal{M}$ is either a hosohedron or a hemi-hosohedron. If $\mathcal{M}$ is a polyhedron, it must be a hosohedron.
Since facets are digons, the monodromy $r_{0}$ commutes with $r_{1}$, and since it commutes with $r_{2}$, it must be in the center of ${\rm Mon}(\mathcal{X})$. This implies that there are at most two vertices, which are the orbits under $\langle r_{1},r_{2}\rangle$ of $\Phi$ and $\Phi^{0}$. If these two orbits are different (which is necessary for $\mathcal{M}$ to be a polyhedron), then $r_{0}$ maps any flag in one orbit to a flag in the other one and $\mathcal{M}$ is a hosohedron. If there is only one vertex, we may rename the flags as $\{1,2,\ldots,2k\}$ in such a way that $r_{1}$ is the permutation $(1,2)(3,4)\cdots(2k-1,2k)$ and $r_{2}$ is the permutation $(2,3)(4,5)\cdots(2k,1)$. Notice that for every $i\in\{1,2,\ldots,2k\}$, one element of $\{r_{1},r_{2}\}$ maps $i$ to $i+1$, while the other one maps $i$ to i1i-1. Since $r_{0}$ commutes with both $r_{1}$ and $r_{2}$ we get that $r_{0}(i+1)=r_{0}(i)\pm 1$. It follows by induction that either $r_{0}(i)=i+j$ for some fixed $j$ and for all $i$ or $r_{0}(i)=j-i$ for some fixed $j$ and for all $i$. Suppose $r_{0}(i)=j-i$ for all $i$. If $j$ is even, then $r_{0}(j/2)=j/2$, a contradiction since $r_{0}$ cannot have fixed points in a maniplex. If $j$ is odd, then $r_{0}((j-1)/2)=(j+1)/2=(j-1)/2+1$, but this is either $r_{1}((j-1)/2)$ or $r_{2}((j-1)/2)$, a contradiction since $r_{0}r_{1}$ and $r_{0}r_{2}$ cannot have fixed points in a maniplex. Therefore, $r_{0}(i)=i+j$ for all $i$ and a fixed $j$. Finally, since $r_{0}$ must have order 2, $j=k$ and $\mathcal{M}$ is a hemi-hosohedron.
Since hosohedra and hemi-hosohedra are all regular, ˜5.2 implies that all maniplexes with only digonal faces are regular. In particular, there are no maniplexes with trivial facet stabilizer with symmetry type graph $\mathcal{X}_{s}$. Therefore ˜5.1 implies the following proposition:
@@
Let $\mathcal{M}$ be a semiregular facet-transitive 3 maniplex with trivial facet stabilizer. Let $\mathcal{X}$ be the symmetry type graph of $\mathcal{M}$. Then, if $\mathcal{X}$ is regular it is isomorphic to a premaniplex in either family 1 or family 2.
@@ 6 Constructing polyhedra with symmetry type graph in family 1
Now we want to assign voltages to the premaniplexes of the previous section to get examples of maniplexes and polytopes with trivial facet stabilizer and with these symmetry type graphs. First, let $\mathcal{X}$ be the premaniplex of family 1 with $n$ sides. We want to find a voltage assignment $\xi$ on $\mathcal{X}$ such that $\mathcal{X}^{\xi}$ has symmetry type $\mathcal{X}$ and $n$-gonal faces. To ensure that $\mathcal{X}^{\xi}$ has $n$-gonal faces it is enough to assign trivial voltage to all the edges of color different than 2. Now let $\alpha_{0},\alpha_{1},\ldots,\alpha_{n-1}$ be the voltages of the edges of color 2 in cyclical order. To make sure that $\mathcal{X}^{\xi}$ is a maniplex we only need to make sure that $\alpha_{i}$ has order exactly 2 for every $i$. And finally, to make sure that $\mathcal{X}^{\xi}$ has symmetry type $\mathcal{X}$, we only need to make sure that no automorphism of $\mathcal{X}$ induces an automorphism of our voltage group $\Gamma=\langle\alpha_{0},\ldots,\alpha_{n-1}\rangle$. Given a base flag $x$ in $\mathcal{X}$, we can can find a closed path $W_{i}$ whose voltage is $\alpha_{i}$ by using edges of colors 0 and 1 to get to the starting point of the dart with voltage $\alpha_{i}$, then using that dart and going back to the base using colors 0 and 1. Given an automorphism $\tau$ of $\mathcal{X}$, the voltage of $W_{i}\tau$ will be the voltage of the image of the dart with voltage $\alpha_{i}$. So, if $\tau$ induces an automorphism of $\Gamma$, this automorphism must be an extension of the permutation that $\tau$ induces on the generators $\{\alpha_{0},\ldots,\alpha_{n-1}\}$. One way to make sure that no non-trivial automorphism of $\mathcal{X}$ induces an automorphism of $\Gamma$ is by using the automorphism group of a regular $n$-polytope $\mathcal{P}$ as $\Gamma$ and making $\alpha_{i}=\rho_{i}$, where $\rho_{i}$ is the distinguished generator of $\Gamma(\mathcal{P})$ that maps a base flag $\Phi$ of $\mathcal{P}$ to its $i$-adjacent flag $\Phi^{i}$. If $\mathcal{P}$ is non-degenerate, then the only possible non-trivial permutation of distinguished generators that can induce an automorphism of $\Gamma(\mathcal{P})$ is the one mapping $\rho_{i}$ to $\rho_{n-1-i}$ for every $i$, as any other permutation would map a pair of non-commuting generators to a pair that commutes. This permutation corresponds to some reflection of $\mathcal{X}$ and it induces an automorphism of the group if and only if $\mathcal{P}$ is self-dual. So, if $\mathcal{P}$ is not self-dual and non-degenerate, $\mathcal{X}^{\xi}$ is a maniplex with symmetry type $\mathcal{X}$. This voltage assignment has been illustrated in Figure˜6, where the label of an edge of the polygon corresponds to the voltage of the darts of color 2 that start on that edge.
@@
[3, Proposition 3.14] Let $(\mathcal{X},\Gamma,\xi)$ be a voltage $(r+1)$-premaniplex. Suppose $\mathcal{X}_{\overline{r}}$ is connected, $\xi(d)=1$ for every dart $d$ of color smaller than $r$ and that $\mathcal{X}^{\xi}$ is a maniplex. Then $\mathcal{X}^{\xi}$ is a polytope if and only if $\mathcal{X}_{\overline{r}}$ is a polytope, and whenever there is a path $W$ with voltage 1 and colors in $[m,r]$, there is also a path $W^{\prime}$ with the same endpoints as $W$ that uses only colors in [m,r1][m,r-1].
In our case $r=2$. Notice that by hypothesis $\mathcal{X}_{\overline{r}}$ is connected, which makes the case $m=0$ trivial. On the other hand, for $m=2$ one only has to notice that there are no edges of color 2 with trivial voltage. Therefore, the only interesting case is when $m=1$. In other words, we want to make sure that if two flags are connected by a path with colors 1 and 2 and voltage 1, then they are either the same flag or 1-adjacent flags.
@@
This condition can be interpreted in terms of the Petrie polygon of $\mathcal{P}$. Typically, a Petrie polygon of a polytope is defined as a cyclical sequence of edges (1-faces) such that whenever $2\leq i\leq n-1$, any $i$ consecutive edges share an $i$-face but any $i+1$ consecutive edges do not. According to this definition some degenerate polytopes and some non-polytopal maniplexes would not have Petrie polygons, so the following alternative is preferred: We will define a Petrie polygon as a flag orbit under the action of the monodromy $\omega_{\pi}:=r_{n-1}r_{n-2}\cdots r_{0}$. The cardinality $\ell$ of a Petrie polygon $P$ is called its length. If we look at the sequence of edges $\{(\omega_{\pi}^{i}\Phi)_{1}\}_{i=0}^{\ell}$ on a non-degenerate polytope, we would see that it satisfies the usual definition of Petrie polygon. Actually, in [24] the sequence $\{(\omega_{\pi}^{i}\Phi)_{1}\}_{i=0}^{\ell}$ is called a Petrie scheme and its sequence of edges is called a Petrie polygon, but in this article we do not make that distinction. Notice that the edge $(\omega_{\pi}^{i}\Phi)_{1}$ is incident to the vertices $(\omega_{\pi}^{i}\Phi)_{0}$ and $(\omega_{\pi}^{i+1}\Phi)_{0}$. In this definition we make sure that every flag is in exactly one Petrie polygon. If any two different flags in $P$ have different vertices (0-faces), we will say that $P$ is simple.
The following lemma will be useful: